OpenStudy (anonymous):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased adult tickets and child tickets.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (ayyookyndall):

Hold on. Im trying to calculate

OpenStudy (anonymous):

kk ^_^

TheSmartOne (thesmartone):

9a+4.5c=72 a+c=11

TheSmartOne (thesmartone):

Solve that ^^ eaution a is adult and c is children

OpenStudy (ayyookyndall):

5 adult tickets 6 child tickets

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/jzdwngtjhk

TheSmartOne (thesmartone):

So yes 5 adults and 6 children..

OpenStudy (ayyookyndall):

5 * 9 = 45 6 * 4.50 = 27 27 + 45 = 72

OpenStudy (anonymous):

the closest i got was 76.5

OpenStudy (anonymous):

thnx

OpenStudy (ayyookyndall):

Oh. Well at least you know the answer now! :) And no problem

TheSmartOne (thesmartone):

and 5+6=11

TheSmartOne (thesmartone):

No problem!